OpenStudy (anonymous):

PLEASE HELP x - 2y = 14 x + 3y = 9 (1, 12) (-1, -12) (12, -1) (12, 1)

OpenStudy (anonymous):

ok all you do is plug them in

OpenStudy (anonymous):

x = 14 + 2y x + 3y = 9 so 14 + 2y + 3y = 9 5y = -5 y = -1 and x + 3(-1) = 9 x = 12

OpenStudy (anonymous):

omg thank you <33333